OpenStudy (anonymous):

Purvi and her friends tipped their waiter a total of ten dollars for dinners that cost nine dollars, eight dollars, eleven dollars and twelve dollars each. Use a percent proportion to solve for the percent tip they left the waiter. Show your work.

OpenStudy (anonymous):

first you will add up the price of all their meals combined: 9+8+7+12=36 then you will make a proportion to find the percent tipped: \[\frac{ 10 }{ 36 }=\frac{ x }{ 100 }\]

OpenStudy (anonymous):

10 for amount tipped. 36 for total meal cost. x for percent tipped because it is unknown. 100 because percent is calculated out of 100. Can you find the percent from here?

OpenStudy (anonymous):

wooops read 'eleven' as 'seven'. the proportion is actually: \[\frac{ 10 }{ 40 }=\frac{ x }{ 100 }\]